1/6x + 1/2 = 2/6, solve x
Please help me solve this equation. THanks

Answer by gonzo(654)

1/6x + 1/2 = 2/6
multiply both sides of the equation by 6x to remove the denominator.
equation becomes
1*6x/6x + 1*6x/2 = 2*6x/6
which becomes
1 + 6x/2 = 12x/6
which becomes
1+3x = 2x
subtracting 2x from both sides of the equation and it becomes
1+x = 0
subtracting 1 from both sides of the equation and it becomes
x = -1
answer is x = -1
-----
another way to solve is as follows:
1/6x + 1/2 = 2/6
subtracting 1/2 from both sides of the equation and it becomes
1/6x = 2/6 - 1/2
right side of equation becomes 2/6 - 3/6 (1/2 is the same as 3/6)
equation then becomes
1/6x = (2-3)/6 = -1/6
multiplying both sides of the equation by 6x gets
6x/6x = 6x*(-1/6)
which becomes
1 = -6x/6
which becomes
1 = -x
which becomes
x = -1 after you multiply both sides of the equation by -1.
answer is x = -1
-----